9. Find the local/absolute extrema off on the closed interval; clearly label your solutions f(x)= x3 – 3x +1 (0,3]
Answer #1

x - 30+ 1 [03] Ques f(n) = Ans) f(n) = 3n²-3. f(x)=0 = 3(021) =0 12) + 1,-1 So, these points are of local manima a minina. S